Suppose that the economy is characterized by the following behavioral equations: C = 180 + 0.70YD 1 = 190 G = 120 T = 100

Please provide the following answers (round response to nearest two decimals):

1. Equilibrium GDP(Y)

2. Disposable Income (YD)

3. Consumption spending (C)
Answer #1

Answer

1)

equilibrium GDP is at

Y=C+I+G

Y=180+0.70Y+190+120

0.3Y=490

Y=1633.33333

The equilibrium GDP is $1633.33333

=====

2)

Disposable income =Y-T=1633.33333-100=1533.33333

=====

3)

Consumption spending is

C=180+0.7*1633.33333

=1323.33333